OpenStudy (beth12345):

Three ants are struggling to move the remains of a bee. Three uncles refuse to help. One ant pushes north with a force of 35mN (millinewton). A second ant pushes east with a force of 12mN. The third ant, an uncooperative little bug, exerts a force which exactly cancels the other two. What is the magnitude of the force exerted by the third ant, and in what direction does she push?

OpenStudy (anonymous):

The forces are at a right angle to each other so you can use, to find the magnitude (the minus sight is to show it's in the opposite, opposing direction: \[r^2 = - \sqrt{x^2 + y^2}\]

OpenStudy (anonymous):

The minus sign is outside the square root. I don't know what happened to the notation of that square root.

OpenStudy (anonymous):

For the direction, use: tan(theta) = x/y I just look at a graph to figure out which of the possible angles it is the angle is.

OpenStudy (beth12345):

thanks

OpenStudy (anonymous):

Welcome